Problem 6:

Ten manufactured parts are randomly selected from a batch where it is believed that the percent nonconforming is 15%. It is important to determine the probability that exactly two out of 10 manufactured parts will be nonconforming. If the random variable X represents the number of nonconforming parts;

  1. What is the probability that exactly two parts are nonconforming?
  1. What is the probability that less than two parts are nonconforming?
Answer #1

(a)
Binomial Distribution

n = 10

p = 0.15

q = 1 -p = 0.85

So,

Answer is:

0.2759

(b)
P(X<2)=P(X=0) + P(X=1)

So,

P(X<2) = 0.5443

So,

Answeris:

0.5443